Compute the following to get the equilibrium price and quantity.


1) Qd= 30-5P


2) Qs= 0+5P



1
Expert's answer
2022-01-19T16:15:49-0500

At equilibrium;

"Q_d=Q_s\\30-5P=0+5P\\30=10P\\P=3"


Substituting P in Qd to get equilibrium quantity ;

"Q=30-5(3)\\
 =15"


Equilibrium price=3

Equilibrium quantity=15